Find the exact value of sin 165°.
Thank You

Answer by fractalier(6550)

Okay we use the addition formula for sine and get
sin 165 = sin(120 + 45) =
sin 120 cos 45 + cos 120 sin 45 =
[sqrt(3)/2 * sqrt(2)/2] + (-1/2)(sqrt(2)/2) =
[sqrt(6) - sqrt(2)] / 4